# Thickness dependent rare earth segregation in magnetron deposited NdCo4.6 thin films studied by Xray reflectivity and Hard Xray photoemission

### v3

@@ III.3.2 Co 2p and Nd 3d spectra
Figure 7 shows the relative atomic concentrations of Co with respect to Nd obtained in the analyzed samples at three diferent probe depths. The Co-to-Nd atomic concentrations were lower than the nominal values in all the cases. In addition, the Co atomic concentration increases linearly with the related of the excited electrons in the 30 nm and 65 nm thick samples. The concentration saturated in the 10 nm thick sample because its thickness was shorter than the related . The linear increase in the concentration with he increasing probe depth agrees with a layer of Nd segregated from the alloy whose thickness has to be a fraction of the used in the experiment. This linear increase was different in the 65 nm film than in the thinner films. This sample has the lowest Co-to-Nd concentration at the lowest , indicating that the segregated layer was the thickest of all the analyzed samples. The estimated thickness of the Nd top layer ranges between 3 and 4 nm, with a Nd atomic concentration in the top layer of approximately four times the concentration in the compound layer, NdCo4.6. This variation in atomic concentration is not far from the obtained by assuming that the top layer is made of Nd2O3, the most common neodymium oxide, which has 2.5 times a neodymium atomic concentration of NdCo4.6. The thicknesses of the layer is in the range of the obtained by XRR and the estimated from the intensity and shape of the photoemission background, as will discussed later. The difference in the slope of the Co-to-Nd atomic concentration versus between the samples with 30 nm and 65 nm thicknesses can only be obtained by assuming different values of for each of the films, which is coherent with the differences in the density of the films at the top layers. This result qualitatively agrees with the differences observed by XRR for the top layer of the two samples which showed a gradation in intensity in the thinner samples and a better defined layer in the thicker sample.
@@ IV CONCLUSIONS
The observed Nd segregation at different stages of growth, together with the progressive increase in PMA energy with thickness, suggests that the origin of PMA in these films is linked to the minimization of strain energy during growth. Specifically, the larger atomic volume of Nd compared to Co imposes strain when Nd incorporates into the Co sublattice. Energy minimization favors Nd atomic environments elongated along the vertical direction, thereby relieving strain and promoting the perpendicular anisotropy observed in these films.

### v4

@@ III.3.2 Co 2p and Nd 3d spectra
Figure 7 shows the relative atomic concentrations of Co with respect to Nd obtained in the analyzed samples at three diferent probe depths. The Co-to-Nd atomic concentrations were lower than the nominal values in all the cases. In addition, the Co atomic concentration increases linearly with the related of the excited electrons in the 30 nm and 65 nm thick samples. The concentration saturated in the 10 nm thick sample because its thickness was shorter than the related . The linear increase in the concentration with the increasing probe depth agrees with a layer of Nd segregated from the alloy whose thickness has to be a fraction of the used in the experiment. This linear increase was different in the 65 nm film than in the thinner films. This sample has the lowest Co-to-Nd concentration at the lowest , indicating that the segregated layer was the thickest of all the analyzed samples. The estimated thickness of the Nd top layer ranges between 3 and 4 nm, with a Nd atomic concentration in the top layer of approximately four times the concentration in the compound layer, NdCo4.6. This variation in atomic concentration is not far from the obtained by assuming that the top layer is made of Nd2O3, the most common neodymium oxide, which has 2.5 times a neodymium atomic concentration of NdCo4.6. The thicknesses of the layer is in the range of the obtained by XRR and the estimated from the intensity and shape of the photoemission background, as will discussed later. The difference in the slope of the Co-to-Nd atomic concentration versus between the samples with 30 nm and 65 nm thicknesses can only be obtained by assuming different values of for each of the films, which is coherent with the differences in the density of the films at the top layers. This result qualitatively agrees with the differences observed by XRR for the top layer of the two samples which showed a gradation in intensity in the thinner samples and a better defined layer in the thicker sample.
@@ IV CONCLUSIONS
The observed Nd segregation at different stages of growth proofs the diffusion of Nd atoms towards the surface of the film during thin film deposition. This fact together with the large compressive strain previously probed by EXAFS and the progressive increase in PMA energy with thickness, suggests that the origin of PMA in these films is linked to the minimization of strain energy during growth. Specifically, the larger atomic volume of Nd compared to Co imposes strain when Nd incorporates into the Co sublattice. Energy minimization favors Nd atomic environments elongated along the vertical direction, thereby relieving strain and promoting the perpendicular anisotropy observed in these films.